OpenStudy (anonymous):

25^(x-2)=125 solve for X

geerky42 (geerky42):

Remember, 25 = 5² and 125 = 5³. Do you know the next step?

OpenStudy (anonymous):

divide 5?

geerky42 (geerky42):

Well, no. Try to get base to be the same so we can cancel it.

OpenStudy (anonymous):

we can multiply one by 3 and the other by 2 so we can get 5^6?

geerky42 (geerky42):

No, it'd mess equation up. Here, let me show you. \[\Large25^{x-2}=125 \Rightarrow (5^2)^{x-2} = 5^3\] See what I did?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

Now, what's the next step?

OpenStudy (anonymous):

do we try to get rid of 5^2?

OpenStudy (anonymous):

or do we use log or ln

geerky42 (geerky42):

Just base. (5)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont get it